The partial probability distribution of X, the number of defective tires on a randomly selected automobile checked at a certain inspection station, is given below. If one of those automobiled is equally likely to have either 2 or 3 defective tires, what is P[(X = 2) U (X = 4)] = P(2 U 4)?

x 0 1 2 3 4
P(x) .54 .12 .20
Answer #1

The sum of all the probabilities should be 1. Hence,

P(0) + P(1) + P(2) + P(3) + P(4) = 1

0.54 + 0.12 + P(2) + P(3) + 0.20 = 1

P(2) + P(3) = 0.14

Since 2 and 3 are equally likely:

P(2) = P(3) = 0.14/2 = 0.07

Therefore,

P[(X = 2) U (X = 4)] = 0.07 + 0.20 = 0.27